OpenStudy (madeleine2):

Write the standard form of the line that passes through the given points. (1, 1) and (3, 4)

OpenStudy (steve816):

Hey there, here you should use the point slope form. Do you know the point slope form?

OpenStudy (steve816):

\[\Huge y-y _{1}=m(x-x _{1})\]

OpenStudy (steve816):

First, you need to find the slope. Slope is \[\Huge \frac{ y _{2}- y _{1}}{ x _{2}- x _{1}}\]

OpenStudy (madeleine2):

(1-3)=m(1-4)?

OpenStudy (madeleine2):

(1-3)/(1-4)=0.666666667

OpenStudy (steve816):

Not quite, you should do (4-1)/(3-1) = m

OpenStudy (madeleine2):

(4-1)/(3-1) = m ok then what

OpenStudy (steve816):

Then, you plug in one of your points to the point slope form!

OpenStudy (steve816):

Let's choose the point (1, 1)

OpenStudy (steve816):

Equation of the point slope form would be:\[\Huge y-1=\frac{ 3 }{ 2 }(x-1)\]

OpenStudy (madeleine2):

but why 3/2 though when it equals -2/-3

OpenStudy (madeleine2):

@steve816